OpenStudy (freemap):

Find the equation of the circle with center at (3, 2) and through the point (5, 4).

OpenStudy (freemap):

@jim_thompson5910

OpenStudy (anonymous):

(x - x1)^2 + ( y - y1)^2 = r^2 where x1 and y1 is the center

OpenStudy (anonymous):

r is the radius of the circle

OpenStudy (anonymous):

you can find the radius using the distance formula between the two points

OpenStudy (freemap):

okay I'm gonna work it out

OpenStudy (freemap):

isn't x the same as x1

jimthompson5910 (jim_thompson5910):

(h,k) is the center r is the radius (x,y) is the point that lies on the circle \[\Large (x-h)^2 + (y-k)^2 = r^2\] \[\Large (x-3)^2 + (y-2)^2 = r^2\] \[\Large (5-3)^2 + (4-2)^2 = r^2\] solve for r

OpenStudy (anonymous):

this way works as well

OpenStudy (freemap):

2²+2²=r² 4+4=8 8x8=64

jimthompson5910 (jim_thompson5910):

you should get to \[\Large r^2 = 8\] but r is NOT equal to 64

jimthompson5910 (jim_thompson5910):

(something)*(something) = 8 what is that "something"?

OpenStudy (freemap):

4x2

jimthompson5910 (jim_thompson5910):

4 times 2 is 8 but I'm asking for two factors which are the same number

jimthompson5910 (jim_thompson5910):

example: 5 times 5 = 25 6 times 6 = 36 etc

OpenStudy (freemap):

i don't know cause 2x2=4 and 4x4=16

jimthompson5910 (jim_thompson5910):

notice how 2*2 = 4 and 3*3 = 9 so whatever that number is, it's between 2 and 3

OpenStudy (freemap):

2.82

jimthompson5910 (jim_thompson5910):

you can apply the square root to both sides to solve for r \[\Large r^2 = 8\] \[\Large \sqrt{r^2} = \sqrt{8}\] \[\Large r = ??\]

jimthompson5910 (jim_thompson5910):

yes roughly 2.828

OpenStudy (freemap):

Thanks,sorry it took so long

OpenStudy (freemap):

sorry it took Me so long

jimthompson5910 (jim_thompson5910):

that's ok